# WildCAT3D: Appearance-Aware Multi-View Diffusion in the Wild

### v1

@@ Abstract
Despite recent advances in sparse novel view synthesis (NVS) applied to object-centric scenes, scene-level NVS remains a challenge. A central issue is the lack of available clean multi-view training data, beyond manually curated datasets with limited diversity, camera variation, or licensing issues. On the other hand, an abundance of diverse and permissively-licensed data exists in the wild, consisting of scenes with varying appearances (illuminations, transient occlusions, etc.) from sources such as tourist photos. To this end, we present WildCAT3D, a framework for generating novel views of scenes learned from diverse 2D scene image data captured in the wild. We unlock training on these data sources by explicitly modeling global appearance conditions in images, extending the state-of-the-art multi-view diffusion paradigm to learn from scene views of varying appearances. Our trained model generalizes to new scenes at inference time, enabling the generation of multiple consistent novel views. WildCAT3D provides state-of-the-art results on single-view NVS in object- and scene-level settings, while training on strictly less data sources than prior methods. Additionally, it enables novel applications by providing global appearance control during generation.
@@ 1 Introduction
We exhaustively compare our method on standard NVS benchmarks and demonstrate superior performance, while training on strictly less data sources than prior works. Our results also show strong performance on diverse scenes captured in tourist photos, including static video generation from single input frames and custom camera trajectories. In addition, our explicit modeling of appearance both allows us to learn from in-the-wild data to successfully generate consistent views of full scenes and also enables novel applications such as interpolation between views of differing appearances or NVS with appearance control via text, as showcased in Figure 1.
@@ 3.2 Appearance-Aware Multi-View Diffusion
Since the appearance embedding used at inference is arbitrary, we can actually use a completely external image to compute the desired appearance embedding and thus perform appearance transfer or appearance-controlled generation. In order to support appearance control through text, we propose to concatenate our model with a text-to-image retrieval model. More specifically, given a text prompt, we use CLIP Radford et al. [2021] to compute its similarity with pre-computed image embeddings from a given database (we use MegaDepth dataset in practice).
@@ 3.4 Implementation Details
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Zero-1-to-3 (released) 6.8720 0.2100 0.5647 128.93 0.0297 10.720 0.2865 0.5258 171.21 0.1126
ZeroNVS (released) 5.7992 0.1113 0.6476 159.96 0.0352 6.9993 0.1240 0.6685 137.04 0.0537
Zero-1-to-3 (MS) 7.6367 0.2755 0.5158 101.94 0.0223 12.921 0.3828 0.4429 67.645 0.0163
ZeroNVS (MS) 8.0191 0.3066 0.4833 87.406 0.0158 13.780 0.4122 0.4057 60.677 0.0139
SD-Inpaint 9.9463 0.3688 0.4951 214.42 0.1067 12.922 0.3996 0.4557 150.11 0.0792
MS NVS 8.7953 0.3930 0.3995 85.959 0.0163 14.056 0.4406 0.3807 64.406 0.0142
WildCAT3D (Ours) 10.766 0.4264 0.3880 57.323 0.0039 14.770 0.4446 0.3522 42.168 0.0050
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Zero-1-to-3 (released) 11.632 0.4384 0.4050 160.20 0.0725 9.0896 0.2413 0.5476 86.892 0.0634
ZeroNVS (released) 9.4869 0.3527 0.4563 123.01 0.0352 7.4711 0.1508 0.6156 69.097 0.0487
Zero-1-to-3 (MS) 14.638 0.5697 0.2722 68.908 0.0024 12.159 0.3665 0.4289 9.7835 0.0023
ZeroNVS (MS) 16.015 0.6304 0.2053 61.117 0.0024 12.900 0.4005 0.3857 9.8382 0.0024
SD-Inpaint 15.541 0.6429 0.2694 118.94 0.0396 12.358 0.3923 0.4245 38.484 0.0242
MS NVS 17.224 0.6661 0.1774 60.013 0.0023 13.397 0.4446 0.3444 11.580 0.0040
WildCAT3D (Ours) 21.579 0.7584 0.1306 24.703 -0.0001 13.916 0.4389 0.3547 9.8705 0.0015
Table 1: Novel-view synthesis benchmarks. We evaluate the single-view setup and compare our results to prior works as reported by Tung et al. [2024]: Zero-1-to-3 [Liu et al., 2023a], ZeroNVS Sargent et al. [2023], a naive inpainting method dubbed SD-Inpaint, and MegaScenes NVS model Tung et al. [2024]). Our method achieves superior performances than baseline methods while using strictly fewer data sources and not requiring aggressive data filtering. Best results are in bold.
@@ 4.2 Additional Applications
In Figures 45, we illustrate additional applications of WildCAT3D. By injecting the appearance embedding of an external image during inference, we may generate novel views of a scene while editing its appearance (e.g. generating a nighttime scene observed from a daytime photo). By concatenating this with CLIP text-to-image retrieval [Radford et al., 2021], we may use text to control the edit (e.g. “sunset”). Finally, by fine-tuning WildCAT3D with two observed input views, we can interpolate between views of a scene with differing appearances to produce a static video with a consistent appearance starting and ending at the two respective poses. At inference time, this uses camera poses along an interpolated trajectory between the two input cameras, along with the appearance of either the start or the end pose injected into each slot.
@@
GT Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
-warp Refer to caption Refer to caption Refer to caption
-warp-app Refer to caption Refer to caption Refer to caption
@@
Figure 6: Qualitative NVS results. (a) We compare predicted views to the ground-truth for MegaScenes NVS model from Tung et al. [2024] (MS NVS) and ours. Our model shows greater consistency with target poses as well as better visual quality, consistent with our quantitative results. (b) Removing warp conditioning (-warp) results in misalignement relative to ground-truth camera poses. Training CAT3D directly on in-the-wild data (i.e. without warps and appearance embeddings, -warp-app) yields inconsistent output images.
@@ 4.4 Ablations
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
WildCAT3D (Ours) 10.766 0.4264 0.3880 57.323 0.0039 14.770 0.4446 0.3522 42.168 0.0050
-warp 9.7950 0.3735 0.4387 54.886 0.0042 13.984 0.4037 0.3948 44.818 0.0056
-warp-app 8.6992 0.2805 0.4913 68.801 0.0069 13.900 0.3899 0.4170 46.800 0.0064
Base CAT3D 10.248 0.3986 0.4238 53.964 0.0042 13.917 0.3991 0.4100 57.955 0.0205
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
WildCAT3D (Ours) 21.579 0.7584 0.1306 24.703 -0.0001 13.916 0.4389 0.3547 9.8705 0.0015
-warp 18.972 0.6698 0.1824 28.709 0.0003 13.453 0.4102 0.3787 9.9337 0.0014
-warp-app 17.728 0.6249 0.2158 31.991 0.0008 12.631 0.3683 0.4221 12.414 0.0026
Base CAT3D 21.893 0.7505 0.1268 21.532 -0.0004 12.910 0.3897 0.4148 19.163 0.0116
Table 2: Quantitative ablation study. We evaluate ablating key parts of our model, namely warp (“warp”) and appearance components (“app”), when trained on in-the-wild data. We also report the performances of the base CAT3D without any in-the-wild finetuning. Best results are in bold.
@@ 5 Conclusion
We have presented the WildCAT3D framework for generalizable novel view synthesis learned from images in-the-wild. By explicitly modeling appearance variations, WildCAT3D unlocks training on abundant and permissively-licensed photo-tourism data, as well as allowing control over the global appearance conditions of generated views. Our results have shown its superior performance on NVS benchmarks and novel applications, while training on strictly less data sources than prior methods.
Limitations and Future Work. Generated views are not guaranteed to be fully consistent, unlike methods using explicit 3D representations. Our architecture could be enhanced by incorporating a video prior, textual conditioning, automatic generation of camera poses, or explicit modeling of transient occlusions. Additional components may help to model semantic variations between images (e.g. holiday decorations) or view-dependent effects such as reflections.
Societal Impact. While scene generation shows promise for positive applications in entertainment and education, we acknowledge the inherent risks of visual generative models to produce disinformation or undesired hallucinations. Our work aims to encourage the adoption of open data for state-of-the-art NVS, while still requiring the same caution in usage as existing generative methods.
@@ A.2 Image Qualitative Results
Figures 8 and 9 illustrate further examples of our outputs on in- and out-of-distribution NVS benchmarks, showcasing our strong performance at predicting novel views from a single observation.
@@ A.3 MegaScenes–Only Ablation
In Table 3 we compare to training end-to-end using data in-the-wild from MegaScenes as the only source of multi-view data (rather than including CO3D and Re10K in multi-view training). While this underperforms our full model including these curated sources of multi-view data, it still achieves competitive performance overall, indicating that MegaScenes alone may be used to learn a strong prior on consistent novel views of scenes despite itself containing inconsistencies between scene views due to appearance variations.
@@
Figure 10: 3D Reconstruction from Generated Views. We show 3D reconstruction results applied to WildCAT3D generations, as described in Section A.4. Reconstructed point clouds and camera positions are shown above, generally matching the expected scene geometry and camera trajectories used for generation. While two generated views are shown for conciseness, 3D reconstruction results are calculated from 15 views.
While the main focus of our work is generating novel 2D views of a scene, we also show that these outputs may be used for downstream 3D reconstruction. In particular, we feed the 15 generated novel views of a scene (using $v=16$ slots for WildCAT3D) into VGGSfM [Wang et al., 2024], a state-of-the-art feed-forward 3D reconstruction pipeline that recovers camera poses and the 3D geometry of the scene. Examples are illustrated in Figure 10, showing that our model’s outputs are sufficiently high-quality and consistent to recover underlying 3D scene geometry. Additionally, the recovered camera trajectories generally match the trajectories used for conditional generation (lateral, circular, and straight trajectories respectively in the examples shown).
@@ B.1 Image Resolution
We train on 512$\times$512 pixel resolution ($64\times 64$ latent resolution); images with other aspect ratios are resized so the longest edge is 512 and padded to be square. Such padding is also used by prior work and evaluation benchmarks (e.g. Gao et al. [2024], Tung et al. [2024]); unlike square cropping, it avoids losing information from the image periphery and allows for generation of images with different aspect ratios in inference. For metric calculations, we resize predictions to 256$\times$256 resolution.

### v2

@@ Abstract
Despite recent advances in sparse novel view synthesis (NVS) applied to object-centric scenes, scene-level NVS remains a challenge. A central issue is the lack of available clean multi-view training data, beyond manually curated datasets with limited diversity, camera variation, or licensing issues. On the other hand, an abundance of diverse and permissively-licensed data exists in the wild, consisting of scenes with varying appearances (illuminations, transient occlusions, etc.) from sources such as tourist photos. To this end, we present WildCAT3D, a framework for generating novel views of scenes learned from diverse 2D scene image data captured in the wild. We unlock training on these data sources by explicitly modeling global appearance conditions in images, extending the state-of-the-art multi-view diffusion paradigm to learn from scene views of varying appearances. Our trained model generalizes to new scenes at inference time, enabling the generation of multiple consistent novel views. WildCAT3D provides state-of-the-art results on single-view NVS in object- and scene-level settings, while training on strictly fewer data sources than prior methods. Additionally, it enables novel applications by providing global appearance control during generation.
@@ 1 Introduction
We exhaustively compare our method on standard NVS benchmarks and demonstrate superior performance, while training on fewer curated data sources than prior works. Importantly, this highlights the strength of our work in leveraging a larger set of samples from existing, permissively licensed imagery captured in-the-wild, rather than relying on heavily curated datasets. Our results also show strong performance on diverse scenes captured in tourist photos, including static video generation from single input frames and custom camera trajectories. In addition, our explicit modeling of appearance both allows us to learn from in-the-wild data to successfully generate consistent views of full scenes and also enables novel applications such as interpolation between views of differing appearances or NVS with appearance control via text, as showcased in Figure˜1.
@@ 3.2 Appearance-Aware Multi-View Diffusion
We note that this derives appearance representations directly from images, rather than using accompanying textual metadata. This has several advantages: Images may not be accompanied by descriptive text, conditioning on text would require additional novel components to bridge text and visual appearance spaces, and important appearance details may be poorly reflected in text, such as precise image aspect ratios and lighting conditions.
Our encoder design follows standard convolutional principles with gradually decreasing spatial dimensions to create a low-dimensional appearance bottleneck. The architecture requires low-dimensional output to serve as an effective information bottleneck. Overall, this uses a minimal network avoiding additional design choices or parameters that would require computationally intensive testing. Further details are provided in the appendix.
@@
Since the appearance embedding used at inference is arbitrary, we can actually use a completely external image to compute the desired appearance embedding and thus perform appearance transfer or appearance-controlled generation. In order to support appearance control through text, we propose to concatenate our model with a text-to-image retrieval model. More specifically, given a text prompt, we use CLIP Radford et al. [2021] to compute its similarity with pre-computed image embeddings from a given database. In practice, we perform this retrieval over approximately 10K images from MegaScenes, covering varying weather and lighting conditions similar to the full dataset.
@@ 3.4 Implementation Details
The additional input channels are handled through a lightweight 1×1 convolution projection layer, making them compatible with the lower input dimension expected by the LDM’s denoising network while adding negligible extra parameters. See the appendix for further details.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$
Zero-1-to-3 (released) 6.872 0.210 0.565 128.9 0.297 10.72 0.287 0.526 171.2 1.126
ZeroNVS (released) 5.799 0.111 0.648 160.0 0.352 6.999 0.124 0.669 137.0 0.537
Zero-1-to-3 (MS) 7.637 0.276 0.516 101.9 0.223 12.92 0.383 0.443 67.65 0.163
ZeroNVS (MS) 8.019 0.307 0.483 87.41 0.158 13.78 0.412 0.406 60.68 0.139
SD-Inpaint 9.946 0.369 0.495 214.4 1.067 12.92 0.400 0.456 150.1 0.792
MS NVS 8.795 0.393 0.400 85.96 0.163 14.06 0.441 0.381 64.41 0.142
WildCAT3D (Ours) 10.77 0.426 0.388 57.32 0.039 14.77 0.445 0.352 42.17 0.050
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$
Zero-1-to-3 (released) 11.63 0.438 0.405 160.2 0.725 9.090 0.241 0.548 86.89 0.634
ZeroNVS (released) 9.487 0.353 0.456 123.0 0.352 7.471 0.151 0.616 69.10 0.487
Zero-1-to-3 (MS) 14.64 0.570 0.272 68.91 0.024 12.16 0.367 0.429 9.784 0.023
ZeroNVS (MS) 16.02 0.630 0.205 61.12 0.024 12.90 0.401 0.386 9.838 0.024
SD-Inpaint 15.54 0.643 0.269 118.9 0.396 12.36 0.392 0.425 38.48 0.242
MS NVS 17.22 0.666 0.177 60.01 0.023 13.40 0.445 0.344 11.58 0.040
WildCAT3D (Ours) 21.58 0.758 0.131 24.70 -0.001 13.92 0.439 0.355 9.871 0.015
Table 1: Novel-view synthesis benchmarks. We evaluate the single-view setup and compare our results to prior works as reported by Tung et al. [2024]: Zero-1-to-3 [Liu et al., 2023a], ZeroNVS Sargent et al. [2023], a naive inpainting method dubbed SD-Inpaint, and MegaScenes NVS model Tung et al. [2024]). Our method achieves superior performance compared to baseline methods while using strictly fewer data sources and not requiring aggressive data filtering. KID* indicates KID values multiplied by ten for readability. Best results are in bold.
@@ 4.2 Additional Applications
In Figures 45, we illustrate additional applications of WildCAT3D. By injecting the appearance embedding of an external image during inference, we may generate novel views of a scene while editing its appearance (e.g. generating a nighttime scene observed from a daytime photo). By concatenating this with CLIP text-to-image retrieval [Radford et al., 2021], we may use text to control the edit (e.g. “sunset”); as seen in the figure, text-based retrieval may effectively find images capturing the desired overall appearance for subsequent injection. Finally, by fine-tuning WildCAT3D with two observed input views, we can interpolate between views of a scene with differing appearances to produce a static video with a consistent appearance starting and ending at the two respective poses. At inference time, this uses camera poses along an interpolated trajectory between the two input cameras, along with the appearance of either the start or the end pose injected into each slot.
@@
GT
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
-warp
Refer to caption Refer to caption Refer to caption
-warp-app
Refer to caption Refer to caption Refer to caption
@@
Figure 6: Qualitative NVS results. (a) We compare predicted views to the ground-truth for MegaScenes NVS model from Tung et al. [2024] (MS NVS) and ours. Our model shows greater consistency with target poses as well as better visual quality, consistent with our quantitative results. (b) Removing warp conditioning (-warp) results in misalignment relative to ground-truth camera poses. Training CAT3D directly on in-the-wild data (i.e. without warps and appearance embeddings, -warp-app) yields inconsistent output images.
@@ 4.4 Ablations
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$
WildCAT3D (Ours) 10.77 0.426 0.388 57.32 0.039 14.77 0.445 0.352 42.17 0.050
-warp 9.795 0.374 0.439 54.89 0.042 13.98 0.404 0.395 44.82 0.056
-warp-app 8.699 0.281 0.491 68.80 0.069 13.90 0.390 0.417 46.80 0.064
Base CAT3D 10.25 0.399 0.424 53.96 0.042 13.92 0.399 0.410 57.96 0.205
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID* $\downarrow$
WildCAT3D (Ours) 21.58 0.758 0.131 24.70 -0.001 13.92 0.439 0.355 9.871 0.015
-warp 18.97 0.670 0.182 28.71 0.003 13.45 0.410 0.379 9.934 0.014
-warp-app 17.73 0.625 0.216 31.99 0.008 12.63 0.368 0.422 12.41 0.026
Base CAT3D 21.89 0.751 0.127 21.53 -0.004 12.91 0.390 0.415 19.16 0.116
Table 2: Quantitative ablation study. We evaluate ablating key parts of our model, namely warp (“warp”) and appearance components (“app”), when trained on in-the-wild data. We also report the performance of the base CAT3D without any in-the-wild finetuning. KID* indicates KID values multiplied by ten for readability. Best results are in bold.
@@ 5 Conclusion
We have presented the WildCAT3D framework for generalizable novel view synthesis learned from images in-the-wild. By explicitly modeling appearance variations, WildCAT3D unlocks training on abundant and permissively-licensed photo-tourism data, as well as allowing control over the global appearance conditions of generated views. Our results have shown its superior performance on NVS benchmarks and novel applications, while training on strictly fewer data sources than prior methods and more fully leveraging existing open web data capturing full scenes. Our experiments, while on a limited scale, demonstrate that modeling appearance enables learning from unfiltered, in-the-wild data, laying a foundation for web-scale NVS training.
Limitations and Future Work. Generated views are not guaranteed to be fully consistent, unlike methods using explicit 3D representations. While our method mitigates degradation due to training on mutually inconsistent data, some visual artifacts persist, such as mild flickering and saturation changes. Our architecture could be enhanced by incorporating a video prior, textual conditioning, automatic generation of camera poses, or explicit modeling of transient occlusions. Additional components may help to model semantic variations between images (e.g. holiday decorations) or view-dependent effects such as reflections. Our method relies on an existing depth estimation pipeline, and errors in predicted depth may propagate to novel views (see appendix). Finally, while our text-based conditioning via image retrieval may provide coarse appearance conditioning, future work could add explicit textual conditioning to allow more fine-grained editing or control over semantic content of scenes.
Societal Impact. While scene generation shows promise for positive applications in entertainment and education, we acknowledge the inherent risks of visual generative models to produce disinformation or undesired hallucinations. The central aim of our work is to encourage the adoption of open data for state-of-the-art NVS, while still requiring the same caution in responsible usage as existing generative methods.
@@ A.2 Image Qualitative Results
Figures˜8 and 9 illustrate further examples of our outputs on in- and out-of-distribution NVS benchmarks, showcasing our strong performance at predicting novel views from a single observation.
@@ A.3 MegaScenes–Only Ablation
In Table˜3 we compare to training end-to-end using data in-the-wild from MegaScenes as the only source of multi-view data (rather than including CO3D and Re10K in multi-view training). While this underperforms our full model including these curated sources of multi-view data, it still achieves competitive performance overall, indicating that MegaScenes alone may be used to learn a strong prior on consistent novel views of scenes despite itself containing inconsistencies between scene views due to appearance variations.
@@
3DGS Novel Rendered Views and Depth Maps
Refer to caption Refer to caption Refer to caption
Figure 10: 3D Reconstruction from Generated Views. We show 3D reconstruction results applied to WildCAT3D generations, as described in Section˜A.4. Reconstructed point clouds and camera positions are shown above, generally matching the expected scene geometry and camera trajectories used for generation. While two generated views are shown for conciseness, 3D reconstruction results are calculated from 15 views. We also show an example of novel views and their normalized depth maps generated using a 3D Gaussian Splatting (3DGS) representation initialized with the sparse reconstruction of the first scene.
While the main focus of our work is generating novel 2D views of a scene, we also show that these outputs may be used for downstream 3D reconstruction. In particular, we feed the 15 generated novel views of a scene (using $v=16$ slots for WildCAT3D) into VGGSfM [Wang et al., 2024], a state-of-the-art feed-forward 3D reconstruction pipeline that recovers camera poses and the 3D geometry of the scene, and use reconstructed sparse scene geometry to initialize a 3D Gaussian Splatting representation Kerbl et al. [2023]. Examples are illustrated in Figure˜10, showing that our model’s outputs are sufficiently high-quality and consistent to recover underlying 3D scene geometry. Additionally, the recovered camera trajectories generally match the trajectories used for conditional generation (lateral, circular, and straight trajectories respectively in the examples shown).
Input Warp Generated View Target View
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Propagation of Depth Estimation Errors. As our method relies on an existing depth estimation module to calculate warps, errors in depth estimation may propagate to novel views. This is illustrated above, as incorrect warps correspond to errors in depth estimation, which are seen to cause resulting novel generated views to deviate from the correct targets.
A.5 Depth Estimation Error Propogation
As our method relies on an existing depth estimation pipeline (monocular depth estimation aligned to a scene’s SfM pointcloud), it may suffer from error propagation when depth is incorrectly estimated. Examples are illustrated in Figure˜11, suggesting that more robust depth estimation may further improve results.
@@ B.1 Image Resolution
We train on $\times512$ pixel resolution ($64\times 64$ latent resolution); images with other aspect ratios are resized so the longest edge is 512 and padded to be square. Such padding is also used by prior work and evaluation benchmarks (e.g. Gao et al. [2024], Tung et al. [2024]); unlike square cropping, it avoids losing information from the image periphery and allows for generation of images with different aspect ratios in inference. For metric calculations, we resize predictions to $\times256$ resolution.